# KeyFace: Expressive Audio-Driven Facial Animation for Long Sequences via KeyFrame Interpolation

### v1

@@ Abstract
Current audio-driven facial animation methods achieve impressive results for short videos but suffer from error accumulation and identity drift when extended to longer durations. Existing methods attempt to mitigate this through external spatial control, increasing long-term consistency but compromising the naturalness of motion. We propose KeyFace, a novel two-stage diffusion-based framework, to address these issues. In the first stage, keyframes are generated at a low frame rate, conditioned on audio input and an identity frame, to capture essential facial expressions and movements over extended periods of time. In the second stage, an interpolation model fills in the gaps between keyframes, ensuring smooth transitions and temporal coherence. To further enhance realism, we incorporate continuous emotion representations and handle a wide range of non-speech vocalizations (NSVs), such as laughter and sighs. We also introduce two new evaluation metrics for assessing lip synchronization and NSV generation. Experimental results show that KeyFace outperforms state-of-the-art methods in generating natural, coherent facial animations over extended durations, successfully encompassing NSVs and continuous emotions.
@@ 1 Introduction
To address these limitations, we propose KeyFace, a novel two-stage approach for generating long and coherent audio-driven facial animations. In the first stage, a keyframe generation model produces a sequence at a low frame rate conditioned on an identity and audio input, spanning multiple seconds and eliminating the need for motion frames. In the second stage, an interpolation model fills in intermediate frames, ensuring smooth transitions and temporal coherence. By dividing generation into two parts, we implicitly separate motion and identity control, resulting in more natural motion and improved identity preservation over time. For longer sequences, this process can be repeated, with the interpolation model generating seamless transitions between segments. In addition to generating realistic, long-term animations, our pipeline allows for emotions that evolve over time, leveraging the keyframe generation model’s broad contextual span.
Our primary contributions can be summarized as follows:
@@
  • New evaluation metrics. We introduce Lipscore and $NSV_{acc}$, two new metrics for evaluating lip synchronization and NSV generation.
@@ 2 Related Works
Audio-driven facial animation methods [29, 55, 70] focus on generating realistic talking head sequences with lip movements synchronised with the target audio input. Early models, such as [55], used GANs, introducing a temporal GAN to generate talking-head videos from a still image and audio input, while Wav2Lip [44] improved lip-sync accuracy with a pre-trained expert discriminator. More recent 3D-aware and head pose-driven methods [7, 70, 72] aimed to capture head motion, though often struggled with artefacts and unnatural movements.
@@
Controllable emotion has recently become a key focus in audio-driven facial animation to create more realistic, empathetic avatars. Most works [19, 17, 37, 10] use discrete emotion labels (e.g. angry or sad) with intensity levels, but this approach lacks expressivity beyond predefined classes. Some approaches use a driving video or audio as a richer emotional source [52, 26, 42, 35], generating a latent representation from the media to drive the animation. However, they require the driving audio or video during inference, limiting expressiveness and restricting explicit control. Continuous emotion conditioning, using valence and arousal, remains underexplored, despite evidence that it better captures emotional complexity [2, 59]. Additionally, few works allow for continuous emotion variation within a video, and those that do are often limited to a small set of emotions [61], likely due to challenges in achieving coherent long-term animation.
@@
Non-speech vocalizations (NSVs), such as laughter and sighs, are essential to human communication [43, 46], providing context beyond the verbal transcript and greatly enhancing the naturalness of casual speech. Despite this, NSVs are often overlooked in audio-driven facial animation, and state-of-the-art models trained only on speech typically perform poorly on NSVs. Recently, two models have aimed to address this gap: Laughing Matters [5], which proposes a diffusion model that can produce realistic laughter videos from still images and audio, and LaughTalk [51], a 3D model that generates both speech and laughter. However, a model capable of handling multiple NSVs in addition to speech has not yet been explored.
@@ 3 Method
Our two-stage approach, outlined in Fig. 2, starts with the generation of temporally distant keyframes. In the second stage, an interpolation model animates the full sequence by filling gaps between the generated keyframes.
@@ 3.2 Keyframe generation
In the first stage, we generate keyframes that capture essential facial expressions and movements corresponding to the audio input over an extended temporal context. These keyframes act as anchor points for the subsequent interpolation stage, ensuring that the final animation accurately reflects both the audio content and the associated emotional expressions. Our architecture is based on Stable Video Diffusion (SVD) [3], a latent diffusion model with a 2D U-Net backbone augmented with 1D temporal layers in the form of residual blocks and attention mechanisms. We generate $T$ keyframes, spaced $S$ frames apart, to capture long-range temporal dependencies efficiently.
@@ 3.4 Audio encoding
  • Timestep Embeddings: We pass $A_{wb}$ through an MLP and add it to the diffusion timestep embeddings $t_{s}\in\mathbb{R}^{C^{s}}$, resulting in $t_{s}^{\prime}=t_{s}+\text{MLP}(A_{wb})$, where $C^{s}$ is the timestep embedding dimension. This further pushes the model to align image and audio frames, improving lip synchronization.
@@ 3.5 Emotion modelling with valence and arousal
To capture complex, continuously changing emotional expressions, we adopt a continuous representation based on valence and arousal. For each frame, we extract valence and arousal using a pre-trained emotion recognition model [47], encode them into sinusoidal embeddings $E_{v},E_{a}\in\mathbb{R}^{C^{s}}$, and add them to the diffusion timestep embedding:
@@ 4.1 Datasets
We train both models on a carefully selected ensemble of datasets. To ensure optimal results, we use only HDTF [71] and a dataset that we collected, comprising 160 hours of speech and 30 hours of NSVs. We also experiment with CelebV-Text [65] and CelebV-HQ [73] but find that excluding these lower-quality datasets benefits training, as we will discuss in Section 4.1. For testing, we use the HDTF test set, and 100 videos randomly selected from CelebV-Text. Additionally, to evaluate our emotion control, we use a test set selected from MEAD [57], as in [52].
@@ 4.2 Evaluation metrics
We evaluate image quality using the aesthetic quality metric from VBench [25], Fréchet Inception Distance (FID) [20], and Learned Perceptual Image Patch Similarity (LPIPS) [69]. For general video quality, we use Fréchet Video Distance (FVD) [54] and the smoothness metric from [25]. We compute the emotion accuracy ($Emo_{acc}$) using the pre-trained emotion recognition model from [47]. We also introduce two new metrics (further details are provided in the supplementary material).
@@
To evaluate the model’s ability to generate NSVs, we train a video classifier to recognize 8 NSV types (”Mhm”, ”Oh”, ”Ah”, coughs, sighs, yawns, throat clears, and laughter) plus speech, for a total of 9 classes. The classifier is based on a pre-trained MViTv2 [36] for video classification on the Kinetics dataset [32]. We then employ it to evaluate the model’s ability to generate the correct NSV type and measure the overall accuracy, denoting this metric as NSV accuracy ($NSV_{acc}$).
@@ 5.1 Quantitative analysis
We present a quantitative comparison against current state-of-the-art methods in Table 1. KeyFace achieves the lowest FID and FVD, indicating higher realism and temporal coherence. Our model also achieves the highest AQ and LPIPS, confirming the visual appeal of our animations. While SadTalker and V-Express achieve the highest smoothness and LipScores, respectively, KeyFace ranks a close second and outperforms both SadTalker and V-Express on the other metrics, demonstrating better performance overall. Figure 3 illustrates FID over time for videos generated by each method, where our two-stage approach maintains consistent quality without degradation. In contrast, Hallo and AniPortrait suffer from significant quality loss over time.
Additionally, the user study results (Elo) show that our model is preferred over other methods, confirming the effectiveness of our approach. A detailed analysis of the results is provided in the supplementary material.
@@ 5.2 Emotional results
To assess our model’s ability to generate accurate emotional expressions, we evaluate it on MEAD [57], which provides discrete emotion labels for each video, and compare it to state-of-the-art emotional talking face generation models in Table 2.
@@ 5.3 Ablation Studies
We present ablation studies on the five main aspects of the model’s performance:
@@
Table 3: Audio encoder ablation on HDTF [71]. For the $NSV_{acc}$, we use the identities from HDTF but change the audio to include NSVs.
@@ 5.4 Qualitative analysis
Figure 8: Results on an out-of-distribution speaker using the same audio. Please refer to our supplementary material for additional video comparisons.
@@ 6 Conclusion
Appendix A Datasets
A.1 Data details
Dataset # Speakers # Videos Duration
Avg. (sec.) Total (hrs.)
HDTF [71] 264 318 139.08 12
CelebV-HQ [73] 3, 668 12, 000 4.00 13
CelebV-Text [65] 9, 109 75, 307 6.38 130
Collected data 824 4, 677 123.15 160
Collected data (NSV) 639 5, 701 18.94 30
Table 8: Overview of the datasets used in the study.
A.2 Preprocessing details
  • Extract 25 fps video and 16 kHz mono audio.
  • Detect and separate scenes using PySceneDetect.
  • Remove clips without active speakers using Light-ASD [38].
  • Estimate landmarks and poses using face-alignment.
  • Crop the video around the facial region across all frames.
Using this pipeline, we curate CelebV-HQ [73] and CelebV-Text [65].
Appendix B Implementation Details
@@
Appendix C Evaluation metrics
C.1 LipScore
@@
C.2 Non-speech vocalization classifier
@@
Our model is trained using a dataset containing video clips of eight different NSV types and speech. The eight NSV classes are: ”Mhm”, ”Oh”, ”Ah”, coughs, sighs, yawns, throat clears, and laughter. During the training process, video clips corresponding to any of these classes are fed into the model. We train using the AdamW optimizer with a learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. The cross-entropy loss is employed as the loss function.
@@
Appendix D User study details
@@ D.1 Additional ablation
D.2 Limitations
@@
D.3 Additional qualitative results

### v2

@@ Abstract
Current audio-driven facial animation methods achieve impressive results for short videos but suffer from error accumulation and identity drift when extended to longer durations. Existing methods attempt to mitigate this through external spatial control, increasing long-term consistency but compromising the naturalness of motion. We propose KeyFace, a novel two-stage diffusion-based framework, to address these issues. In the first stage, keyframes are generated at a low frame rate, conditioned on audio input and an identity frame, to capture essential facial expressions and movements over extended periods of time. In the second stage, an interpolation model fills in the gaps between keyframes, ensuring smooth transitions and temporal coherence. To further enhance realism, we incorporate continuous emotion representations and handle a wide range of non-speech vocalizations (NSVs), such as laughter and sighs. We also introduce two evaluation metrics for assessing lip synchronization and NSV generation. Experimental results show that KeyFace outperforms state-of-the-art methods in generating natural, coherent facial animations over extended durations, successfully encompassing NSVs and continuous emotions. See the project page for visualizations and code.
@@ 1 Introduction
To address these limitations, inspired by keyframe-based approaches [67, 69] initially introduced by [43], we propose KeyFace, a novel two-stage approach for generating long and coherent audio-driven facial animations. In the first stage, a keyframe generation model produces a sequence at a low frame rate conditioned on an identity and audio input, spanning multiple seconds and eliminating the need for motion frames. In the second stage, an interpolation model fills in intermediate frames, ensuring smooth transitions and temporal coherence. By dividing generation into two parts, we implicitly separate motion and identity control, resulting in more natural motion and improved identity preservation over time. For longer sequences, this process can be repeated, with the interpolation model generating seamless transitions between segments. In addition to generating realistic, long-term animations, our pipeline allows for emotions that evolve over time, leveraging the keyframe generation model’s broad contextual span.
Our main contributions can be summarized as follows:
@@ 2 Related Works
Audio-driven facial animation methods [30, 58, 75] generate realistic talking-head sequences with audio-synchronized lip movements. Early models, such as [58], used GANs, introducing a temporal GAN to generate talking-head videos from a still image and audio input, while Wav2Lip [46] improved lip-sync accuracy with a pre-trained expert discriminator. More recent 3D-aware and head pose-driven methods [7, 75, 77] aimed to capture head motion, though often struggled with artefacts and unnatural movements.
@@
Controllable emotion has recently become a key focus in audio-driven facial animation to create more realistic, empathetic avatars. Most works [19, 17, 38, 10] use discrete emotion labels (e.g. angry or sad) with intensity levels, but this approach lacks expressivity beyond predefined classes. Some approaches use a driving video or audio as a richer emotional source [55, 27, 44, 36, 49], generating a latent representation from the media to drive the animation. This latent representation can sometimes be interpolated to control the resulting emotion [26]. However, they require the driving audio or video during inference, limiting expressiveness and restricting explicit control. Continuous emotion conditioning, using valence and arousal, remains underexplored, despite evidence that it better captures emotional complexity [2, 62]. Additionally, few works allow for continuous emotion variation within a video, and those that do are often limited to a small set of emotions [64], likely due to challenges in achieving coherent long-term animation.
@@
Non-speech vocalizations (NSVs), such as laughter and sighs, significantly enhance human communication [45, 48] by providing context beyond words and increasing speech naturalness. Despite this, NSVs are often overlooked in audio-driven facial animation, and state-of-the-art models trained only on speech typically perform poorly on NSVs. Recently, two models have aimed to address this gap: Laughing Matters [5], which proposes a diffusion model that can produce realistic laughter videos from still images and audio, and LaughTalk [54], a 3D model that generates both speech and laughter. However, a model capable of handling multiple NSVs in addition to speech has not yet been explored.
@@ 3 Method
Our two-stage approach, outlined in Fig. 2, starts with the generation of temporally distant keyframes. In the second stage, an interpolation model animates the full sequence by filling gaps between the generated keyframes. Our architecture builds upon Stable Video Diffusion (SVD) [3], with further architectural details and key distinctions provided in Appendix B.
@@ 3.2 Keyframe generation
In the first stage, we generate keyframes that capture essential facial expressions and movements, guided by the audio over an extended temporal context. These keyframes act as anchor points for the subsequent interpolation stage, ensuring that the final animation accurately reflects both the audio content and the associated emotional expressions. We generate $T$ keyframes, spaced $S$ frames apart, to capture long-range temporal dependencies efficiently.
@@ 3.4 Audio encoding
  • Timestep Embeddings: We pass $A_{wb}$ through an MLP and add it to the diffusion timestep embeddings $t_{s}\in\mathbb{R}^{C^{s}}$, yielding $t_{s}^{\prime}=t_{s}+\text{MLP}(A_{wb})$, where $C^{s}$ is the timestep embedding dimension. This encourages alignment between image and audio frames.
@@ 3.5 Emotion modelling with valence and arousal
To capture complex, continuously changing emotional expressions, we adopt a continuous representation based on valence and arousal. For each frame, we extract valence and arousal using a pre-trained emotion recognition model [50], encode them into sinusoidal embeddings $E_{v},E_{a}\in\mathbb{R}^{C^{s}}$, and add them to the diffusion timestep embedding along with the audio embeddings:
@@ 4.1 Datasets
We train both models on HDTF [76] and a dataset that we collected, comprising 160 hours of speech and 30 hours of NSVs. We also experiment with CelebV-Text [70] and CelebV-HQ [78] but find that excluding these lower-quality datasets benefits training. For testing, we use the HDTF test set, and 100 videos randomly selected from CelebV-Text. Additionally, to evaluate our emotion control, we use a test set selected from MEAD [60], as in [55].
@@ 4.2 Evaluation metrics
We evaluate image quality using the aesthetic quality metric from VBench [25], Fréchet Inception Distance (FID) [20], and Learned Perceptual Image Patch Similarity (LPIPS) [74]. For general video quality, we use Fréchet Video Distance (FVD) [57] and the smoothness metric from [25]. We compute the emotion accuracy ($Emo_{acc}$) using the pre-trained emotion recognition model from [50]. We also introduce two new metrics, further details are provided in Appendix D.
@@
To evaluate the model’s ability to generate NSVs, we train a video classifier to recognize 8 NSV types ("Mhm", "Oh", "Ah", coughs, sighs, yawns, throat clears, and laughter) plus speech, for a total of 9 classes. The classifier is based on a pre-trained MViTv2 [37] for video classification on the Kinetics dataset [33]. We then employ it to evaluate the model’s ability to generate the correct NSV type and measure the overall accuracy, denoting this metric as NSV accuracy ($NSV_{acc}$).
@@ 5.1 Quantitative analysis
We present a quantitative comparison against current state-of-the-art methods in Table 1. KeyFace achieves the lowest FID and FVD, indicating higher realism and temporal coherence. Our model also achieves the highest AQ and LPIPS, confirming the visual appeal of our animations. While SadTalker and V-Express achieve the highest smoothness and LipScores, respectively, KeyFace ranks a close second and outperforms both SadTalker and V-Express on the other metrics, demonstrating better performance overall. Figure 3 illustrates FID over time for videos generated by each method, where our two-stage approach maintains consistent quality without degradation. In contrast, Hallo and AniPortrait suffer from significant quality loss over time. To ensure fairness in evaluation, we also report results for a variant of our model trained exclusively on HDTF in Appendix F.
Additionally, the user study results (Elo) show that our model is preferred over other methods, confirming the effectiveness of our approach. A detailed analysis of the results is provided in Appendix E.
@@ 5.2 Emotional results
To assess our model’s ability to generate accurate emotional expressions, we evaluate it on MEAD [60], comparing it to state-of-the-art models in Table 2.
@@ 5.3 Ablation Studies
Table 3: Audio encoder ablation on HDTF [76]. For $NSV_{acc}$, we use HDTF identities with audio containing NSVs.
@@ 5.4 Qualitative analysis
Figure 8: Results on out-of-distribution id using the same audio. Please refer to our project page for additional video comparisons.
@@ 6 Conclusion
Supplementary Material
Appendix A Model Details
Implementation
@@ Appendix A Model Details
Inference speed
One limitation of our model is that it does not yet support real-time generation. Nevertheless, our two-stage approach is faster than competing diffusion-based models, particularly because it allows batching, unlike autoregressive methods. We present an inference speed comparison (Table 9), measured in seconds per frame. Real-time inference could potentially be achieved through distillation methods (e.g., UFOGen), which we leave for future work.
V-Express [59] Hallo [65] AniPortrait [63] EchoMimic [11] Keyface
3.36 1.9 0.44 0.76 0.26
Table 9: Seconds per frame comparison for baseline models.
Appendix B Comparison with SVD
Our method builds upon Stable Video Diffusion (SVD) [3] by introducing carefully designed architectural and task-specific adaptations. These modifications distinctly set our approach apart from prior work. We highlight the primary differences below.
Audio Conditioning
While SVD primarily conditions on the initial frame to predict subsequent video frames, our method extends this capability by conditioning on both an identity frame and audio inputs to drive video generation. To the best of our knowledge, we are the first to employ conditioning based on outputs from two distinct audio encoders (WavLM [8] and BEATs [9], allowing simultaneous processing of speech and non-speech audio.
Emotional Conditioning
Unlike the original SVD architecture, our approach incorporates additional control over emotional expression. We demonstrate that training emotional models exclusively with pseudo-labels for valence and arousal achieves robust and consistent performance.
Loss Functions
SVD employs only the EDM loss [31]. In contrast, we use two additional pixel-space losses along with a weighted loss that specifically targets the lower region of generated images.
Guidance
Whereas SVD solely employs vanilla classifier-free guidance (CFG) [21], we provide an in-depth investigation into optimal guidance techniques tailored specifically to each stage of our pipeline. We found that, for the keyframe model, assigning different CFG weights to identity and audio conditions leads to better performance and improved robustness compared to classical CFG. Additionally, since interpolation requires greater flexibility in head movement, we employed autoguidance [32] to dynamically balance guidance, resulting in enhanced overall video quality.
Appendix C Datasets
C.1 Data details
Dataset # Speakers # Videos Duration
Avg. (sec.) Total (hrs.)
HDTF [76] 264 318 139.08 12
CelebV-HQ [78] 3, 668 12, 000 4.00 13
CelebV-Text [70] 9, 109 75, 307 6.38 130
Collected data 824 4, 677 123.15 160
Collected data (NSV) 639 5, 701 18.94 30
Table 10: Overview of the datasets used in the study.
C.2 Preprocessing details
  • Extract 25 fps video and 16 kHz mono audio.
  • Detect and separate scenes using PySceneDetect.
  • Remove clips without active speakers using Light-ASD [39].
  • Estimate landmarks and poses using face-alignment.
  • Crop the video around the facial region across all frames.
Using this pipeline, we curate CelebV-HQ [78] and CelebV-Text [70].
However, even after filtering the datasets, we found that many samples contain editing effects and/or occlusions that are not detected. Examples include visible hands, camera movement, editing effects, and occlusions, which we found occur in 20 % of videos even after our cleaning process, as illustred in Figure 9. Since these artefacts don’t correlate with speech, they can’t be replicated by the model, hindering performance as shown in Section 5.3.
Figure 9: Illustration of bad examples in CelebV-HQ [78] and CelebV-Text [70].
Appendix D Evaluation metrics
D.1 LipScore
@@
WER on unseen datasets
We additionally evaluate our state-of-the-art lipreader [41] on HDTF and find that it achieves a 21 % WER, demonstrating strong performance on unseen data and further supporting LipScore’s validity.
D.2 Non-speech vocalization classifier
@@
Our model is trained using a dataset containing video clips of eight different NSV types and speech. The eight NSV classes are: "Mhm", "Oh", "Ah", coughs, sighs, yawns, throat clears, and laughter. During the training process, video clips corresponding to any of these classes are fed into the model. We train using the AdamW optimizer with a learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. The cross-entropy loss is employed as the loss function.
@@
NSVs performance boundaries
To demonstrate and understand the effectiveness of $NSV_{acc}$ across individual NSVs, we present a confusion matrix on the validation set of the data used to train $NSV_{acc}$ (Fig.15, left). Although the model achieves good overall performance, certain NSVs are frequently confused, such as “Oh” with “Ah,” “Sigh” with “Mhm,” and “Yawn” with “Cough.”
Additionally, we demonstrate that our model can generate visually distinct NSVs (Fig.15, right) with few confusions by generating 10 videos per NSV category and speech.
Figure 15: NSV confusion matrix for generated (left) and validation (right) videos.
Appendix E User study details
@@ Appendix F Additional ablation
Training on HDTF only
To ensure a fair comparison with baseline models, we retrain our model exclusively on publicly available data (i.e. HDTF [76]), removing all non-public sources. Although this leads to a decrease in performance, our model still outperforms baseline methods trained on larger datasets. We emphasize that most existing methods rely on private datasets; therefore, to maintain fairness, we curated our dataset to have comparable scale in terms of total hours and number of speakers as described in Section C.1.
Method FID $\downarrow$ FVD $\downarrow$ LipScore $\uparrow$
KeyFace (HDTF only) 19.49 165.06 0.28
Table 12: Results of pipeline trained on HDTF only.
Appendix G Limitations
@@
Appendix H Additional qualitative results